# Feedforward equilibrium trajectory optimization with GSPulse

### v1

@@ 1.1 Motivation
On many tokamaks, portions of the pulse routinely use only feedforward actuation when feedback control is difficult or unavailable. For example, the sequence for plasma initiation usually includes driving the coils along preprogrammed trajectories to apply large voltages for creating plasma breakdown conditions and sustaining plasma burnthrough. This is often done in feedforward-only before transitioning to feedback control. Some reasons for this are that during the highly dynamic early portion of the pulse the phase delay associated with feedback control might be unacceptably high, and that, at low plasma current the signal-to-noise ratio is lower and some real-time reconstruction algorithms may not converge reliably [1] (although simpler non-Grad-Shafranov algorithms could suffice).
@@ 1.2 Pulse simulating vs pulse planning
![[Uncaptioned image]](x2.png)
An additional difference is that, due to the convergence strategy and inherent vertical stabilization (discussed further in section 2.1), the GSPulse pulse planner can take larger time steps than a pulse simulator while remaining vertically stable. This allows for a reduced computational cost of the solution. Some notable examples of forward evolutive pulse simulators are DINA [2], FGE [3], FreeGSNKE [4], and GSevolve [5].
@@ 1.3 Comparison against previous approaches
Several earlier solvers have also treated the inverse-evolutive optimization problem, including CEDRES++ [6] and an algorithm by Blum [7], which have now been incorporated into the NICE code [8]. These solvers use Sequential Quadratic Programming (SQP) methods to optimize for sequences of equilibria performing a similar task as GSPulse. The main differences of these solvers versus GSPulse are: the exact formulation of the cost functional where GSPulse includes additional terms on the derivatives and smoothness of trajectories, and that the SQP-based solvers iterate on the full state vector solution directly, whereas GSPulse splits the iteration into optimizing the vacuum conductor evolution (via Model Predictive Control formulation) and Picard iterations of the plasma state. This framework allows GSPulse to re-use sensitivity matrices used in the optimization, and forgo additional gradient or adjoint state calculations. The implicit assumption is that it is sufficient to simply update the effect of the plasma contribution to the cost function at each iteration, and not necessary to use the gradient of the cost w.r.t plasma contribution within the optimization step. This structure is very beneficial for computational efficiency although perhaps reduces the optimality of the final solution.
Some other codes within this design space are Tokamaker [9], which also includes some time-dependent effects and can solve for sequences of equilibria, but without performing global trajectory optimization. Additionally, recent work on the FBT [10] equilibrium code has extended it to the dynamic case to create the FBT-Evolutive pulse planner, by incorporating the GSPulse equations presented in the next sections (i.e., MPC-like global vacuum optimization coupled with Grad-Shafranov Picard iterations).
@@ 2 Algorithm principles
In this equation, is the cost, is the time period of interest, is the shaping coil power supply voltages, and is the vector of isoflux shaping errors which can include signals like coil currents, shape control point flux errors, field at the x-points, and the boundary flux or loop voltage. are weighting matrices for each cost term. In the implementation, the cost function is discretized to discrete time points and equilibria. Note that there are quadratic penalties on the signal values, as well as on the signal first and second derivatives (with respect to time). First-derivative weights are used to penalize trajectories that ramp too aggressively, and second-derivate weights are used to penalize non-smooth trajectories. Using principles derived from Model Predictive Control (MPC), the cost function subject to the FBEE constraints can be reformulated into the form of a standard constrained quadratic program (QP). The standard form QP can be solved by standard software packages such as MATLAB’s quadprog.
@@
The FBEE governing equations are (using a COCOS-17 coordinate system [11]):
See section A.1 for the definitions of each variable. The first equation is the well-known Grad-Shafranov equation [12, 13] for axisymmetric equilibrium force balance, the second describes axisymmetric circuit dynamics within a tokamak, and the last describes the plasma current evolution. The plasma current governing equation is sometimes presented in different forms. The form here used by GSPulse is the scalar Ejima equation [14] which describes how the plasma surface voltage must compensate for plasma resistance and changes to the plasma current and inductance. Note that Ejima equation (and GSPulse) do not consider non-inductive current, but could be extended be replacing the resistive voltage term above with [15]. GSPulse can also be configured to accept the loop voltage or boundary flux directly, instead of computing it from the Ejima model.
The employed solution strategy is to use a Picard iteration scheme for the Grad-Shafranov plasma flux distribution and couple this with a dynamic optimization. This is analogous to the Picard iteration scheme employed in many static equilibrium solvers, that alternates between updating the conductor currents and the Grad-Shafranov equilibrium. The major difference is that in the GSPulse case we include the dynamics as a constraint in the optimization and also solve for all of the voltages and currents over the time window of interest. Figure 1 gives a cartoon of the convergence strategy used in GSPulse, where we alternate between the global MPC-like optimization and plasma Picard updates of the flux distribution.
@@
Configuration: In this step the user defines a standard set of inputs which includes the tokamak geometry, mutual inductance matrices, coil and vessel resistances, core plasma properties needed to solve the Grad-Shafranov equation such as the P’ and FF’ profile shapes, constraints on the plasma properties such as the the plasma current, thermal energy, or internal inductance, and the target shapes and weights. Note that GSPulse itself does not derive mutual inductance and Green’s function tables from the tokamak geometry, but instead has been configured to accept geometry and inductance tables computed from both the TokSys [16] and MEQ [17] suites of equilibrium codes.
For the inductive plasma current drive model, the user can either specify a waveform for plasma resistance, which is used in the Ejima equation, or directly specify the target boundary flux evolution. Shape targets can be created with the shape editor GUI (fig. 3). Weights are specified for various shaping parameters and coil current and coil voltage characteristics. These are used to tune the solution to give a desired trajectory – for example they can be used to tune how important achieving a particular strike point is, or to penalize non-smooth trajectories for the currents.
@@
Grad-Shafranov Picard iteration: Given the conductor currents and profiles from the iteration, perform a Picard update of the Grad-Shafranov equation to find the flux distribution at the iteration. GSPulse has the option of performing this step natively, or via calling subroutines of the MEQ suite [18] of equilibrium codes.
@@ 2.1 Vertical stabilization during convergence iterations
is unstable when applied repeatedly without feedback [19]. However, this tends to be more of a problem for reconstruction solvers and forward solvers. For example, RT-EFIT [20] and the original LIUQE [18] explicitly add a vertical shift free parameter that could be updated across iterations, and forward simulators like FreeGSNKE [4], FGE [3], and NICE forward mode [8] avoid Picard iterations in favor of Newton or Newton-like methods.
However, as an inverse solver, GSPulse tends to be robust against this vertical instability in practice. This is linked to the shape target objective within the optimization, which applies corrective action between iterations to keep the plasma shape from drifting. The stability is also related to the time resolution of the GSPulse configuration. When very small timesteps are used (roughly on the order of the vertical instability timescale), the enforced dynamic constraint prevents GSPulse from taking sufficiently large stabilizing action and the Picard vertical instability can reappear. With larger timesteps, GSPulse is generally stable given appropriate weighting of the cost function objectives. The ability to take large time steps is beneficial for computational time and makes GSPulse an efficient tool for exploring design space.
@@ 3.1 Validation against NSTX-U 204660
GSPulse has been validated against experimental results and simulations for several tokamaks including NSTX-U [21] and MAST-U [22]. Figure 4 illustrates a validation study against NSTX-U shot 204660. The objective of this GSPulse run is to re-create the exact progression of equilibria and coil currents that was achieved during the actual feedback-controlled pulse [23], as defined by the NSTX-U EFIT01 reconstruction [24]. Since GSPulse internally enforces the FBEE model, we expect that any modeling discrepancies or implementation errors would prevent the algorithm from being able to exactly match both the target shapes and coil currents, up to the limit of accuracy for EFIT01.
@@
The initial condition (equilibrium, coil currents, and vessel currents) are provided from an EFIT [25] reconstruction at ms. The evolution of the boundary flux is specified as a direct target based on the EFIT01 (magnetics-only) loop voltage. Core plasma properties (stored thermal energy, , and P’ and FF’ profile basis function shapes) are also obtained from the EFIT01 values.
@@
One of the emergent features from this run, observable in the equilibria of fig. 4, is that the outer strikepoint moves radially outwards throughout the pulse, even though the inner strike point position, x-point position, and core boundary shape remain relatively fixed. This NSTX-U pulse used a subset of the divertor shaping coils and consequently could not control both the outer strike point and x-point positions. The motion of the strike point is driven by a change in the time-varying current of the OH coil which does not supply a uniform field, a known feature of the NSTX-U design [26]. Since GSPulse is adequately capable of modeling this type of effect, it could be a useful operator tool for evaluating the tradeoffs in shape evolution for different choices of coils and coil limits under time-varying conditions.
@@ 3.3 Design of SPARC strikesweep scenario and validation against FGE
We also validate GSPulse against the FGE [3] code during a strikepoint sweeping simulation for the SPARC tokamak [27]. FGE has been used and tested extensively on a number of tokamaks. It was also the simulation environment for a reinforcement learning magnetic controller [28] that succesfully controlled the TCV tokamak.
@@
The FGE feedback simulation is configured to use the same initial condition as GSPulse. Since FGE is a pulse simulator, we need to specify feedback control within the loop. A vertical controller is designed that uses the VSC coils to apply feedback on the plasma vertical position. Additionally, a PID radial controller is designed that uses the PF4U/L coils to apply feedback on the plasma radial outer gap. Additionally, the voltages computed by GSPulse are used as feedforward inputs to the simulation. In summary, the voltage commands applied during the simulation are a combination of vertical feedback, radial feedback, and GSPulse-computed voltages.
@@
The pulse-planning (GSPulse) and pulse-simulating (FGE) results are shown in figs. 6 and 7. There is strong agreement between these models, again illustrating that GSPulse is reliably solving the FBEE problem and can be used predictively. As seen in fig. 7, there is a little bit of drift in the PF4 current which also has a small impact on the inner strike point position. This drift is a result of using the PF4 coils for outer gap feedback control in addition to feedforward voltage actuation, and could be compensated by also adding a command to track the feedforward currents. Smaller differences could also be achieved by designing the feedback simulation to more closely match the feedforward-only case by removing derivative control (which adds some noise to the voltage) and using the GSPulse-achieved shapes as shape targets, however, the simulation as depicted is more representative of how feedforward targets would be added to the actual SPARC feedback controller.

### v2

@@ 1.1 Motivation
On many tokamaks, portions of the pulse routinely use only feedforward actuation when feedback control is difficult or unavailable. For example, the sequence for plasma initiation usually includes driving the coils along preprogrammed trajectories to apply large voltages for creating plasma breakdown conditions and sustaining plasma burnthrough. This is often done in feedforward-only before transitioning to feedback control. Some reasons for this are that during the highly dynamic early portion of the pulse the phase delay associated with feedback control might be unacceptably high, and that, at low plasma current the signal-to-noise ratio is lower and some real-time reconstruction algorithms may not converge reliably Wai et al. [2022] (although simpler non-Grad-Shafranov algorithms could suffice).
@@ 1.2 Pulse simulating vs pulse planning
An additional difference is that, due to the convergence strategy and inherent vertical stabilization (discussed further in section 2.1), the GSPulse pulse planner can take larger time steps than a pulse simulator while remaining vertically stable. This allows for a reduced computational cost of the solution. Some notable examples of forward evolutive pulse simulators are DINA Khayrutdinov and Lukash [1993], FGE Carpanese [2021], FreeGSNKE Amorisco et al. [2024], and GSevolve Welander et al. [2019].
@@ 1.3 Comparison against previous approaches
Several earlier solvers have also treated the inverse-evolutive optimization problem, including CEDRES++ Heumann et al. [2015] and an algorithm by Blum Blum et al. [2019], which have now been incorporated into the NICE code Faugeras [2020]. These solvers use Sequential Quadratic Programming (SQP) methods to optimize for sequences of equilibria performing a similar task as GSPulse. The main differences of these solvers versus GSPulse are: the exact formulation of the cost functional where GSPulse includes additional terms on the derivatives and smoothness of trajectories, and that the SQP-based solvers iterate on the full state vector solution directly, whereas GSPulse splits the iteration into optimizing the vacuum conductor evolution and Picard iterations of the plasma state. This framework allows GSPulse to re-use sensitivity matrices used in the optimization, and forgo additional gradient or adjoint state calculations. The implicit assumption is that it is sufficient to simply update the effect of the plasma contribution to the cost function at each iteration, and not necessary to use the gradient of the cost w.r.t plasma contribution within the optimization step. This structure is very beneficial for computational efficiency although perhaps reduces the optimality of the final solution.
Another code within this design space is Tokamaker Hansen et al. [2024], which also includes some time-dependent effects and can solve for sequences of equilibria, but without performing global trajectory optimization. Additionally, the FBT Hofmann [1988] equilibrium code has recently been extended to treat the dynamic case.
@@ 2 Algorithm principles
In this equation, is the cost, is the time period of interest, is the shaping coil power supply voltages, and is the vector of isoflux shaping errors which can include signals like coil currents, shape control point flux errors, field at the x-points, and the boundary flux or loop voltage. are weighting matrices for each cost term. In the implementation, the cost function is discretized to discrete time points and equilibria. Note that there are quadratic penalties on the signal values, as well as on the signal first and second derivatives (with respect to time). First-derivative weights are used to penalize trajectories that ramp too aggressively, and second-derivate weights are used to penalize non-smooth trajectories.
The cost function subject to the FBEE governing equations can be reformulated into a standard-form quadratic program (QP) that can be solved by software packages such as MATLAB’s quadprog. This is done by building a prediction model to determine the relationship between actuator commands at every time step and their impact on output errors at every time step. The prediction model technique is similar in principle to the technique used in Model Predictive Control (MPC). Note that unlike traditional MPC, GSPulse performs open-loop optimization; it does not operate in a simulation-type environment where a different QP is solved at every time step. An additional difference is the nonlinear terms of the FBEE model are not linearized, but instead treated exactly and updated at each iteration.
@@
The FBEE governing equations are (using a COCOS-17 coordinate system Sauter and Medvedev [2013]):
See section A.1 for the definitions of each variable. The first equation is the well-known Grad-Shafranov equation Grad and Rubin [1958], Shafranov [1966] for axisymmetric equilibrium force balance, the second describes axisymmetric circuit dynamics within a tokamak, and the last describes the plasma current evolution. The plasma current governing equation is sometimes presented in different forms. The form here used by GSPulse is the scalar Ejima equation Ejima et al. [1982] which describes how the plasma surface voltage must compensate for plasma resistance and changes to the plasma current and inductance. Note that Ejima equation (and GSPulse) do not consider non-inductive current, but could be extended be replacing the resistive voltage term above with Romero [2010]. GSPulse can also be configured to accept the loop voltage or boundary flux directly, instead of computing it from the Ejima model.
The employed solution strategy is to use a Picard iteration scheme for the Grad-Shafranov plasma flux distribution and couple this with a dynamic optimization. This is analogous to the Picard iteration scheme employed in many static equilibrium solvers, that alternates between updating the conductor currents and the Grad-Shafranov equilibrium. The major difference is that in the GSPulse case we include the dynamics as a constraint in the optimization and also solve for all of the voltages and currents over the time window of interest. Figure 1 gives a cartoon of the convergence strategy used in GSPulse, where we alternate between the open-loop trajectory optimization and plasma Picard updates of the flux distribution.
@@
Configuration: In this step the user defines a standard set of inputs which includes the tokamak geometry, mutual inductance matrices, coil and vessel resistances, core plasma properties needed to solve the Grad-Shafranov equation such as the P’ and FF’ profile shapes, constraints on the plasma properties such as the the plasma current, thermal energy, or internal inductance, and the target shapes and weights. Note that GSPulse itself does not derive mutual inductance and Green’s function tables from the tokamak geometry, but instead has been configured to accept geometry and inductance tables computed from both the TokSys Humphreys et al. [2007] and MEQ epf [2025] suites of equilibrium codes.
For the inductive plasma current drive model, the user can either specify a waveform for plasma resistance, which is used in the Ejima equation, or directly specify the target boundary flux evolution. When using Ejima mode, the plasma internal inductance is computed via numerical integration of the equilibrium and updated each iteration. Shape targets can be created with the shape editor GUI (fig. 3). Weights are specified for various shaping parameters and coil current and coil voltage characteristics. These are used to tune the solution to give a desired trajectory – for example they can be used to tune how important achieving a particular strike point is, or to penalize non-smooth trajectories for the currents.
@@
Grad-Shafranov Picard iteration: Given the conductor currents and profiles from the iteration, perform a Picard update of the Grad-Shafranov equation to find the flux distribution at the iteration. GSPulse has the option of performing this step natively, or via calling subroutines of the MEQ suite Moret et al. [2015] of equilibrium codes.
@@ 2.1 Vertical stabilization during convergence iterations
with a free-boundary edge condition is unstable when applied repeatedly without feedback Johnson et al. [1979]. However, this tends to be more of a problem for reconstruction solvers and forward solvers. For example, RT-EFIT Ferron et al. [1998] and the original LIUQE Moret et al. [2015] explicitly add a vertical shift free parameter that could be updated across iterations, and forward simulators like FreeGSNKE Amorisco et al. [2024], FGE Carpanese [2021], and NICE forward mode Faugeras [2020] avoid Picard iterations in favor of Newton or Newton-like methods.
However, as an inverse solver, GSPulse tends to be robust against this vertical instability. This is linked to the user setting a shape target objective within the optimization, which applies corrective action between iterations to keep the plasma shape from drifting. The stability is also related to the time resolution of the GSPulse configuration. Smaller time steps tend to be more difficult to stabilize. The reasoning is that the numerical (unphysical) instability of the equilibrium Picard iteration does not depend on the time resolution, while it must be counteracted by currents in the coils and conducting structures. These do depend on the time resolution, as the circuit dynamic model (eq. 3) constrains the currents such that , that is, the amount of compensating current that can be provided decreases with the time step size. When very small time steps are used, at or below the vertical instability timescale, then the vertical instability can reappear and convergence can be sensitive to the exact weights and tuning parameters.
This is illustrated in fig. 4. In this example, we use GSPulse to design 100ms of a flattop Ip=8.7MA pulse for SPARC (Creely et al. [2020]) and maintain an up-down symmetric double null equilibrium. The vertical instabilty growth rate of this equilibrium is computed via MEQ to be Hz (ms). The time resolution for this pulse is varied from 2-20ms (between 6 and 51 equilibria each) and the maximum value of the magnetic axis Z-position across the sequence of equilibria is plotted. The largest time step of 20ms has exponential convergence in driving the z-axis to exactly zero. At intermediate time steps, the sequence converges to a small sub-mm fixed value, while the 2ms and 5ms resolution pulses go vertically unstable. The exact conditions for whether the vertical stability is developed or avoided depend on interactions between the plasma equilibrium, the time resolution, and the shaping tuning weights, and a detailed study is beyond the scope of this article. However, we note that in typical pulse planning applications it is often desired to evaluate larger time steps over long time horizons, conditions over which the algorithm is more robust against vertical instability leading to few issues in practice.
2.2 Weight design and specification
An important element of achieving a specific pulse trajectory is configuring the weights used by the optimizer. GSPulse has been implemented to optimize for a number of signals related to voltage, current, and isoflux error measurements, that is, the flux or field at specified locations or linear combinations thereof. Isoflux-based signals are used, in contrast to spatial-based signals (like the shape control gaps), because this results in linearizations that are not equilibrium-dependent. It is discussed in more detail in A.2 (step 3) that there is a computational advantage to this approach which allows for large matrix calculations to be re-used.
The error signals currently supported are: power supply voltage, coil currents, linear combinations of coil currents (for example, to penalize up-down asymmetry in coil currents), the absolute flux at specified locations, the relative flux at specified locations (that is, the difference between flux measured at point A vs point B, which can be used for example to ensure that the flux at specified shape control points is equal to the flux at a desired touch point or x-point), and the radial and vertical field at specified locations (for example, to form an x-point by driving the field at a specified location to zero).
The cost associated with any single error signal, at a single time step is:
A useful heuristic is to design the weights such that cost for each error signal is of order unity. This is achieved by specifying the weights such that they are inversely proportional to the square of typical error range values:
To illustrate, consider the following typical pulse scenario objective (the example in section 3.5 approximately matches this): the plasma begins limited, and then transitions to a lower single null, diverted phase. The shape should be controlled relatively accurately, meaning that large scale features like plasma volume, elongation, and triangularity should be actualized. However, it is also desired to minimize coil voltages and currents and therefore some tradeoff in shape is allowed to reduce actuator usage.
To achieve this, one could design the weights as follows (the description of these weights are modified from the section 3.5 example, which has a slightly different objective):
To control the shape while it is limited, we assign weights to penalize the difference in flux between the shape control points and the touch point location. When it is diverted, we penalize the difference between shape control points and the target x-point location. The weights are specified as a function of time and we smoothly vary between these sets of weights at the limited-diverted transition time.
On SPARC, a 1cm spatial shape control gap error corresponds to 0.01Wb (near the x-points) up to 0.4Wb (at the outer midplane). Applying the inverse-square heuristic this corresponds to a weight of . For simplicity, we assign a weight of to all control points.
To form the x-points, we assign a target of at the desired locations. We would like to form these with approximately 1cm precision, which for an 8.7MA SPARC equilibrium corresponds to T so the weight is taken as . The time-varying weight is assigned such that it is zero during most of the limited phase but transitions so that the x-point is formed slightly before the plasma diverts.
To form a lower single null equilibrium we control the relative flux difference between the upper and lower x-points and assign it a non-zero target value. Since 1cm flux difference at the outer midplane corresponds to 0.4Wb, the target flux difference is assigned such that . The weight is assigned equal to the other control point flux weights.
To control the surface voltage, we penalize the time derivative of the absolute value of the flux at the plasma boundary because . From the Ejima equation we have that and on SPARC is approximately unity. Taking the desired precision to be kA/s at MA results in a typical error of Wb/s, resulting in a weight of .
The inverse-square heuristic is also applied to the coil currents and voltages, resulting in weights for these parameters of and . Additionally, to find smooth trajectories, we apply a weight on the second time derivative of the coil currents. The weight is initialized at a small value and increased as much as possible until it is observed to affect the shape no more than several mm, with a final value of .
@@ 3.1 Validation against NSTX-U 204660
GSPulse has been validated against experimental results and simulations for several tokamaks including NSTX-U Berkery et al. [2024] and MAST-U Harrison et al. [2019]. Figure 5 illustrates a validation study against NSTX-U shot 204660. The objective of this GSPulse run is to re-create the exact progression of equilibria and coil currents that was achieved during the actual feedback-controlled pulse Boyer et al. [2018], as defined by the NSTX-U EFIT01 reconstruction Sabbagh et al. [2001]. Since GSPulse internally enforces the FBEE model, we expect that any modeling discrepancies or implementation errors would prevent the algorithm from being able to exactly match both the target shapes and coil currents, up to the limit of accuracy for EFIT01.
@@
The initial condition (equilibrium, coil currents, and vessel currents) are provided from an EFIT Lao et al. [1985] reconstruction at ms. The evolution of the boundary flux is specified as a direct target based on the EFIT01 (magnetics-only) loop voltage. Core plasma properties (stored thermal energy, , and P’ and FF’ profile basis function shapes) are also obtained from the EFIT01 values.
@@
One of the emergent features from this run, observable in the equilibria of fig. 5, is that the outer strike point moves radially outwards throughout the pulse, even though the inner strike point position, x-point position, and core boundary shape remain relatively fixed. This NSTX-U pulse used a subset of the divertor shaping coils and consequently could not control both the outer strike point and x-point positions. The motion of the strike point is driven by a change in the time-varying current of the OH coil which does not supply a uniform field, a known feature of the NSTX-U design Menard et al. [2012]. Since GSPulse is adequately capable of modeling this type of effect, it could be a useful operator tool for evaluating the tradeoffs in shape evolution for different choices of coils and coil limits under time-varying conditions.
@@ 3.3 Design of SPARC strikesweep scenario and validation against FGE
We also validate GSPulse against the FGE Carpanese [2021] code during a strike point sweeping simulation for the SPARC tokamak. FGE has been used and tested extensively on a number of tokamaks. It was also the simulation environment for a reinforcement learning magnetic controller Degrave et al. [2022] that controlled the TCV tokamak, demonstrating FGE’s validity for predictive performance.
@@
The FGE feedback simulation is configured to use the same initial condition as GSPulse. Since FGE is a pulse simulator, we need to specify vertical feedback control within the loop. A feedback controller is designed using the VSC (Vertical Stability Coil) as the primary actuator. The results of the GSPulse computation are used as feedforward inputs to the simulation, so that the output voltages supplied to FGE at each simulation time step are a sum of the shape feedforward and vertical feedback controller commands.
@@
The pulse-planning (GSPulse) and pulse-simulating (FGE) results are shown in figs. 7 and 8. There is strong agreement between these models, again illustrating that GSPulse is reliably solving the FBEE problem and can be used predictively. Figure 8 shows time traces for the strike point locations and representative currents and voltages for 2 of the 19 coils. GSPulse and FGE agree on the strike point positions to within 4mm at all times (note that because of the strike point angle-of-incidence this generally corresponds to mm precision of the plasma boundary). One interesting point is that the target strikesweep was specified as a triangle wave, but the GSPulse-computed voltage trajectories (for example the DV1L voltage trajectory) can differ significantly from a triangle wave, suggesting a sophisticated input-output relationship that could be difficult to design by hand.